Tyrone is twice as old as his daughter. In six years Tyrone's age will be three times what his daughters age was six years ago. How old are they at present?
Answer by longjonsilver(2297)

Let t = age now of Tyrone
Let d = age now of daughter

t = 2d --eqn1

Second piece of info is (t+6) = 3(d-6) --eqn2

eqn2 becomes t+6 = 3d-18. Sub into it, eqn1, to get 2d+6 = 3d-18.
2d+6 = 3d-18
+6 = d-18
d-18 = 6
d = 24
and from eqn1, t=2x24 --> t=48


jon.